In the formation of acid rain, sulfur dioxide reacts with oxygen and water in the air to form sulfuric acid. Write the balanced chemical equation for the reaction. If 9.82 g SO2, react with excess oxygen and water, how many grams of H2SO4 are produced?
1
Expert's answer
2020-04-29T13:28:26-0400

Answer:

The balanced reaction equation is:

2SO2 + O2 + 2H2O "\rightarrow" 2H2SO4

As you can see, the number of the moles of SO2 reacted is equal to the number of the moles of H2SO4 produced:

"n(SO_2) = n(H_2SO_4)"

The molar mass "M" of SO2 is 64.06 g/mol. Let's calculate the number of the moles of SO2:

"n(SO_2) = \frac{m}{M} = \frac{9.82}{64.06} = 0.153" mol

The molar mass "M" of H2SO4 is 98.08 g/mol. Then, the mass of H2SO4 produced is:

"m(H_2SO_4) = n\u00b7M = 0.153\u00b798.08 =15.04" g